OpenStudy (lostinlife15):

What is the equation of a line that passes through (-2, 3) and (1, -6)? A. Y=x+5 B. Y=-3x+9 C. Y=-3x-3 D. Y=-1/3x+7/3

OpenStudy (studygurl14):

Two methods: Use slope formula to find teh slope and then a point to find b or plug in both points into answer choices to see what works

OpenStudy (kittiwitti1):

Slope formula: given points\[(x_{1},y_{1})\text{ and }(x_{2},y_{2})\]use\[m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (studygurl14):

slope-intercept form y = mx + b Once you find m (slope), then plug in one of the given points for x and y to find b

OpenStudy (kittiwitti1):

The answer choices are in the slope-intercept format,\[y=mx+b\]So basically your answer choices are A) m = 1 B) m =-3 C) m = -3 D) m = \(-\frac{1}{3}\) Your next step would be to apply the Point-Slope formula (below) to calculate the complete equation.\[y-y_{1}=m(x-x_{1})\]After you isolate the \(y\) you should get the slope-intercept form.

OpenStudy (lostinlife15):

Alright thank you!